You are here: Foswiki>Commissioning Web>Com20080518 (19 May 2008, AndrewRakich)Edit Attach

Commissioning Notes from 17-18 May 2008 MST

Observers: Doug M on mountain and Andrew R in Tucson.

T.O.: Randy G.

Summary

Highlight: Achieved "closed loop" active optics ("closed" by Doug pushing the "send" button in his IDL routine) with good convergence on the terms we were targetting (z4, z5, z6, z7, z8 and z11). In 1.5" seeing coefficients converged and held at under 50 nm per term. Some refinements still required for on-axis before we go off-axis. Numerous elevation faults and as well several azimuth tracking faults caused by enclosure. Details below.

Nighttime Activities

4:45 Clouds pass and we are open

5:00 Can not communicate to the IRTC. Power cycled the IRTC in the treehouse, but still no communication. John say we have to reboot the Windows computer

5:17 Communication to the IRTC is established. Temp set to 243 and peltier turned on.

5:20 Could not see the dim star on the IRTC

5:21 Decide to slew to a brighter star, but now the IRTC is not able to communicate to the CSQ(?). Get error that PresetTelescope failed and Invalid buffer:RPCERROR. Restarted irs three times. Randy restartes PCS and CSQ. No communication. Call John.

5:30 RPCERROR means one of the TCS subsystems the is needed to point the telescope is not working properly. Randy checked and OSS was not running. Started that and we could pointto a star.

5:30 pointto WT10_249

5:45 Clouds move in. Can sometimes see a star

5:55 Homed all AGW stages. No problems.

6:30 regained guiding

7:05 checking focus coupling into Z11

7:06 Polaris monitor down, estimate <1" seeing from WFS images

7:08 lost guiding elevation axis failure

7:14 reacquired guiding

7:17 lost elevation drive again

7:20 regained tracking

7:21 lost el axis again

7:30 after some wire jiggling in el axis power supply regained tracking

7:34 Polaris 1.3"

7:35 acquired WFS image

7:39 Predicted that a sign error on the z4 coefficient would give -2.50 microns of z4 when we sent 500 nm of Z11. Measured -2.50 microns of z4, seems to confirm sign error on new z4 coefficient in PSF

7:53 closed ActO loop by hand, worked pretty well, taking new block 5 images with near zero aberrations WFSC000025-29

7:56 taking new block with +1000 z6 WFSC000030-34

8:03 el axis failure

8:13 failed to acquire guidestar when it was ~100 pixels from bottom of image, failed twice, bumped it up with pt modify IE and successfully acquired

8:28 GCSL died unexpectedly

8:50 attempting closed loop ActO using Z4, Z5, Z6, Z7, Z8 and Z11, coefficients as follows: (z4, 0.8), (z5, +1), (z6, -1), (z7, -1.5), (z8, 1.7) and (z11, 1.5). when sending z11 we are sending additional z4 by a factor of -5X the requested z11 term, to counter the sign error in PSF.

8:52 after correcting a small bug try again , and lose el axis

8:56 try again, sequence starting at wfsc000044 sent corrections through to wfsc000051, 52 not sent, unwrapping rotator. Got good convergence for a first try!

9:18 new star, 70 deg EL. starting "closed loop" ActO starting wfsc000053, discovered a sign error in z11 (another one) fixed, "closed loop" starting at wfsc000055-60

9:35 1.5" seeing. Deliberately driven in bad aberration: +1.2 microns z4. z5, 6, 7 and 11 = +600 nm, z8 = -600 nm. guider image looks like **. Will try to correct with "closed loop" ActO.

9:47 still frigging around with trying to get guiding, multiple el axis failures

9:52 1.5" seeing. Deliberately driven in bad aberration: -300 microns z4. z5, 6, 7 and 11 = +300 nm, z8 = -600 nm. guider image still looks like **. Will try to correct with "closed loop" ActO. Image sequence begins wfsc000062-67. In moderately bad seeing everything converged to under 50 nm bar focus that bnounced around a little.

10:04 wfsc000069-73 first of a sequence of 5 "zero" aberration WFS images

10:07 driving in +500 z9. taking new sequence of images starting from wfsc000074-78

10:11 another 5 zero driven aberrations wfsc000079-83

10:17 momentarily lost el, starting again with wfsc80-84 (aberration free sequence 2) killed as we need to slew to new star.

10:40 Recovered tracking after Az failure caused by enclosure. Randy issues issuetrak 1361. Starting sequence of 5 zero aberration reference images, GCS dies

10:46 restart GCS,

11:04 hit x limit, manually inputting some global offsets to give us more room: delx = -0.47 mm, Ry = -10".

11:06 el drive failure

11:15 running a couple of loops of ActO to set zero for sequence of 5 reference WFSC images starting at WFSC000089-93

11:18 sequence of 5 with +500 z10 wfsc000094-98

11:23 sequence of 5 back to zero forced aberrations wfsc000099-103

11:28 z22 driven to 300 nm, taking 5 images starting at WFSC0000104-108

11:30 doing a couple of ActO iterations to clean up image, then some through focus images on the guide camera and IRTC.

11:36 WFS is starting to wash out before we achieved convergance, so we'll start again on this tomorrow

Waiting....

-- DougMiller and AR- 18 May 2008
Topic revision: r4 - 19 May 2008, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback